News

Network status: Spectral camera repaired; Tenerife telescope malfunction

On Monday, 19 October, from 17 - 21 UT, LCO's observatory portal will be closed while we address some software maintenance tasks. You will not be able to submit new observing requests during this time, and the portal website will show an error message. 
The cable to the Spectral camera shutter on the 2m telescope on Haleakala has been replaced, and the camera (fs02) is again accepting requests for science observations. We will be monitoring the camera to verify that the shutter continues to function normally.
We've been having problems getting the 0.4-meter telescope on Tenerife (camera kb29) to move, so it's been taken out-of-service. Our TelOps team is working with the site staff at Teide Observatory to diagnose the problem.
Another Sinistro camera is ready to be installed! The latest Sinistro, fl11, has arrived in Australia, and we expect that it will be installed on the 1m telescope in Dome B next week. After installation,  the telescope will be unavailable for science observations while the camera undergoes the commissioning procedure.
Share on Twitter Share on Facebook